# Scattering Amplitudes and Conservative Binary Dynamics at ${\cal O}(G^{5})$ without Self-Force Truncation

### v1

@@ Abstract
We compute the complete potential-graviton contributions to the conservative radial action and scattering angle for two non-spinning bodies in general relativity, accurate through fifth order in Newton’s constant and including second-order self-force (2SF) effects. The calculation is carried out in the scattering-amplitude framework, combining the double copy, effective field theory, and multi-loop integration techniques based on integration by parts and differential equations. To address a major computational bottleneck, we develop improved integration-by-parts algorithms that render calculations at this order tractable. The amplitude is presented as a series expansion, following the strategy used earlier in maximal supergravity. For the first self-force sector, which involves only polylogarithmic functions, we also provide a closed-form analytic expression. For the second self-force sector, as in earlier supergravity work, we find nontrivial cancellations among contributions related to integrals supported on Calabi–Yau geometry.
@@
Basic setup— We adopt the setup used to compute the potential-mode 2SF contributions in maximal supergravity Bern et al. (2025a), which in turn parallels the three-loop, amplitude-based framework developed in Refs. Bern et al. (2021, 2022). In Einstein gravity, the main qualitative difference is a substantial increase in the complexity of the integral reductions: in the absence of supersymmetry the relevant integrands typically carry higher tensor rank, leading to significantly larger IBP systems.
@@
The classical limit corresponds to large angular momentum, $J\gg\hbar$, or equivalently to small momentum transfer $q=p_{1}+p_{4}$ (with all momenta taken incoming). At loop level, this selects regions where loop momenta $\ell$ scale as $\ell\sim|q|\ll m_{i}\sim\sqrt{s}$, so that, within the EFT-based method of regions Beneke and Smirnov (1998), the classical contributions arise from the soft region. At order ${\cal O}(G^{5})$ this yields a classical amplitude ${\cal M}^{\text{cl}}_{5}\propto q^{2}\ln(-q^{2})$, corresponding in position space to a conservative $1/r^{5}$ potential. To assist with the expansion in the soft region and the subsequent evaluation of integrals we follow Refs. Landshoff and Polkinghorne (1969); Parra-Martinez et al. (2020) and employ special variables where all mass and momentum-transfer dependence factorizes from the integrals, leaving nontrivial functions of the single dimensionless boost parameter $\sigma={p_{1}\cdot p_{2}}/{(m_{1}\,m_{2})}={1}/{\sqrt{1-v^{2}}}$, or equivalently the relative velocity $v$.
In this Letter, we focus on the potential region, i.e. the subset of the soft region describing instantaneous interactions between massive particles, where loop momenta scale as $\ell=(\ell^{0},\bm{\ell})\sim|q|(v,\mathbf{1})$ with $v\ll 1$. Classical observables are then extracted from the radial action, which is related to the massive-scalar $2\to 2$ scattering amplitude by Bern et al. (2021),
@@
Computational Setup— We extract classical conservative 5PM observables from the classical scattering amplitude of massive particles at $\mathcal{O}(G^{5})$ in general relativity, i.e. at four-loop order. We construct the integrand by first obtaining the requisite tree amplitudes via the double-copy, sewing them into generalized cuts and which we then expand in the soft region through ${\cal O}(|\bm{q}|^{2})$. The resulting expressions are then mapped to a cut-integrand basis Bern et al. (2025b) to remove redundant labeling and subsequently merged into a soft-expanded integrand by picking each term exactly once while verifying cut consistency.
@@
The IBP reduction is challenging due to the proliferation of irreducible numerators, doubled propagators, intricate topologies, and the large set of master integrals. We carry out the reductions with FIRE7 Smirnov and Zeng (2025), leveraging improved finite-field performance von Manteuffel and Schabinger (2015); Peraro (2016). In addition, we refine the choice of “seed integrals,” extending our previous strategy Bern et al. (2024b) and related developments Driesse et al. (2024); Guan et al. (2025). (see Kira 3 Lange et al. (2025) for a public implementation and Refs. Larsen and Zhang (2016); Wu et al. (2023); Abreu et al. (2017, 2018) for analogous seeding schemes in earlier non-Laporta approaches.)
@@
As in Ref. Bern et al. (2025a), rather than seeking closed-form expressions in terms of special functions, in the potential region, we exploit that the master integrals are meromorphic in $\bar{\sigma}\equiv\sigma-1$ near $\bar{\sigma}=0$ and thus admit a local Laurent expansion about the static limit. Solving the differential equations by the Frobenius method yields a series solution,
@@
with coefficients $\vec{c}_{n}(\epsilon)$ that are rational functions of the dimensional regulator $\epsilon$. The lower bound $n_{0}$ is determined by the leading velocity scaling in the static limit, while $n_{\max}$ is chosen to reach the target accuracy, here $\mathcal{O}(\bar{\sigma}^{15})$. Expanding $\hat{M}(\sigma,\epsilon)$ around $\bar{\sigma}=0$, we solve the system recursively in $\bar{\sigma}$: at each order the unknowns follow from linear systems whose $\epsilon$ dependence is reconstructed Wang (1981); Wang et al. (1982) from numerical evaluations over prime fields. This approach builds on earlier applications in collider physics Mistlberger (2018); Moriello (2020); Pozzorini and Remiddi (2006); Hidding (2021). Finally, after planarizing the most challenging 2SF classical contributions Bern et al. (2024b, 2025a) using the support of cut matter lines Saotome and Akhoury (2013); Akhoury et al. (2021), we obtain 371 boundary conditions, fixed from the static limit of scalar integrals that contain both classically singular and regular terms. Of these, only 95 enter the genuinely classical observables.
@@
5PM Results— Once expressed in terms of boundary integrals, the results organize themselves into well-defined eikonal sums Saotome and Akhoury (2013); Akhoury et al. (2021). The diagrams within a given eikonal sum are related by permuting vertices on matter lines, resulting in a sum over eikonal propagators that gives an on-shell $\delta$-function, forcing them to be cut. Fig. 2 gives the planar representatives defining the eikonal sums relevant for the classical amplitude, with the remaining ones given by by permuting the attachment point of the potential-region graviton lines, represented by dashed lines, on the matter lines. See Refs. Bern et al. (2024b, 2025a) for further details on eikonal sums in the 5PM amplitudes.
@@
We solved the system of differential equations through $\mathcal{O}(\bar{\sigma}^{15})$; however, spurious singularities limit the result to the first eleven orders in the velocity expansion. Including the overall factor of $G^{5}$, this corresponds to terms through 10PN order.
@@
where for the sake of brevity we wrote explicitly only terms through 7PN order; further terms through $\mathcal{O}(\bar{\sigma}^{6})$ (10PN order) are included in the ancillary file GR_potential_radial_action.m. Such divergent terms in the potential-graviton contribution to the radial action are expected to be related to the odd-in-velocity part of the energy loss at lower (in this case 4PM) order Bini and Damour (2017); Bini et al. (2020); Blanchet et al. (2020),
@@
Terms through $\mathcal{O}(\bar{\sigma}^{6})$ (10PN) are included in the ancillary file GR_potential_radial_action.m. The ancillary file also contains the exact analytic expressions for the 0SF and 1SF contributions, where all special functions but polylogarithms cancel, as observed earlier in Ref. Driesse et al. (2024). We note here the presence of terms proportional to $\pi^{2}$, which are absent from the analogous results in ${\cal N}=8$ supergravity Bern et al. (2025a). In our exact 1SF potential radial action, as well as in the 1SF conservative results of Driesse et al. (2024), such terms originate in (multiple) polylogarithmic functions. This is unlike the 4PM case, where $\pi^{2}$ terms tag only elliptic integrals. It will be interesting to determine the special functions that come with $\pi^{2}$ factors at 2SF order.
@@
Checks— Our result passes several nontrivial checks. First, as pointed out above, certain spurious higher-order poles in the dimensional regularization parameter $\epsilon$ cancel in the radial action and the remaining divergence matches the energy loss at 4PM order. At 0SF we reproduce the exact probe limit result presented in e.g. Ref. Kol et al. (2022). Following the discussion in Ref. Bern et al. (2019a), the leading in velocity terms of the scattering angle are predicted by lower-order iterations. These velocity iterations arise from the integral representation of the scattering angle in terms of the radial momentum, see Eq. (2), and they determine the angle at orders $\bar{\sigma}^{-j}$, $j=2,3,4,5$,
@@
where $A=\frac{55}{6}{-}16\ln 2$. We checked using the 4PM angle $\chi_{4}$ Bern et al. (2021) that our results in Eq. (17) agree with this prediction. Note that the ${1}/{\epsilon}$ divergence of the 2SF angle is partially checked by the velocity iteration consistency at order $\bar{\sigma}^{-2}$ and the presence of a similar divergence of the potential-region 4PM angle $\chi_{4}$. Notably, a combination of all these checks targets all classical eikonal sums, reinforcing the correctness of our result. It would also be interesting to compare with the scattering angle that follows from the PN Hamiltonian of Refs. Blümlein et al. (2020, 2021, 2022).
Discussion— We include in Fig. 3 plots of the finite part of the radial action that illustrate some of their features. The left panel illustrates the relative importance of the different self-force components in the finite part of the radial action as a function of velocity. The right panel indicates that for $v\gtrsim.7$, higher-order terms and resummations are required to improve convergence. Once the radiation-mode contributions are incorporated, it will be essential to obtain accurate representations of the series solutions across the entire physical region. Several complementary strategies are available. As emphasized in Ref. Bern et al. (2025a), the choice of expansion variables plays a critical role in convergence. The conventional PN expansion in the velocity $v$ is not optimal, as it can introduce limitations from singularities outside the physical domain. A more suitable choice of expansion variable is $x$, related to the boost variable by $\sigma=(1+x^{2})/(2x)$, which leads to better convergence properties in the physical region $x\in[0,1]$. This strategy is incorporated in Fig. 3.
@@
Conclusions— In this Letter we computed the complete potential-mode contributions to the amplitude and conservative scattering angle at 5PM order. Our calculations show that the computationally demanding 2SF component can be brought under systematic control. We employed state-of-the-art Feynman integration tools, as published by two of the authors Smirnov and Zeng (2025), along with various private enhancements. Our framework closely parallels that employed in the corresponding analysis of maximal supergravity Bern et al. (2025a), augmented by substantial improvements in IBP seeding strategies, which should likewise transfer to multi-loop collider-physics applications.
After obtaining the four-scalar integrand and performing the IBP reduction, we solved the differential equations for all master integrals as a series expansion about the static limit $\bar{\sigma}\to 0$. The resulting solution determines the classical amplitude (radial action) at $\mathcal{O}(G^{5})$ through 10PN order. Extending the series to substantially higher orders will be useful once radiation contributions are included. It would be interesting to promote the series solution to an exact analytic result, with full control over generalized polylogarithms, elliptic integrals, and Calabi–Yau integrals.
Radiation modes must also be included. While observables are independent of the precise split between conservative and dissipative dynamics, a definite scheme is required for comparison with other approaches Wheeler and Feynman (1949); Damour and Esposito-Farese (1996); Damour (2016); Barack and Pound (2019); Driesse et al. (2024). For waveform modeling, the PM results can be embedded in the EOB framework Khalil et al. (2022); Buonanno et al. (2024), potentially supplemented by resummations exploiting the singularity structure of observables Damour and Rettegno (2023); Long et al. (2024) or information from the high energy limit Alessio et al. (2025). Finally, extending the results to bound orbits requires an analytic continuation using appropriate prescriptions, as discussed in Refs. Khalil et al. (2022); Bini and Damour (2024); Dlapa et al. (2024, 2025). Such continuations are nontrivial at 4PM and beyond due to the nonlocal-in-time contributions from the tail effect. Once the complete 5PM results are available, it would be very interesting to compare them with high-precision numerical relativity Long et al. (2025) or self-force studies Barack et al. (2023) to assess the magnitude of the effects of time nonlocality (which may be small cf. e.g. Ref. Buonanno et al. (2024)) and the size of the still-to-be-calculated 6PM contributions as well as to identify improved resummations of the PM series Long et al. (2024).

### v2

@@ Abstract
We compute the complete potential-graviton contributions to the conservative radial action and scattering angle for two non-spinning bodies in general relativity, accurate through fifth order in Newton’s constant and including second-order self-force (2SF) effects. The calculation is carried out in the scattering-amplitude framework, combining the double copy, effective field theory, and multi-loop integration techniques based on integration by parts and differential equations. To address a major computational bottleneck, we develop improved integration-by-parts algorithms that render calculations at this order tractable. The post-Minkowskian amplitude is presented as a series expansion, following the strategy used earlier in maximal supergravity. For the first self-force sector, which involves only polylogarithmic functions, we also provide a closed-form analytic expression. For the second self-force sector, as in earlier supergravity work, we find nontrivial cancellations among contributions related to integrals supported on Calabi–Yau geometry.
@@
Basic setup— We adopt the setup used to compute the potential-mode 2SF contributions in maximal supergravity Bern et al. (2025a), which in turn parallels the three-loop, amplitude-based framework developed in Refs. Bern et al. (2021, 2022). In Einstein gravity, the main qualitative difference is a substantial increase in the complexity of the integral reductions: in the absence of supersymmetry, the relevant integrands typically carry higher tensor rank, leading to significantly larger IBP systems.
@@
The classical limit corresponds to large angular momentum, $J\gg\hbar$, or equivalently to small momentum transfer $q=p_{1}+p_{4}$ (with all momenta taken incoming). At loop level, this selects regions where loop momenta $\ell$ scale as $\ell\sim|q|\ll m_{i}\sim\sqrt{s}$, so that, within the EFT-based method of regions Beneke and Smirnov (1998), the classical contributions arise from the soft region. At order ${\cal O}(G^{5})$ this yields a classical amplitude ${\cal M}^{\text{cl}}_{5}\propto q^{2}\ln(-q^{2})$, corresponding in position space to a conservative $1/r^{5}$ potential. To assist with the expansion in the soft region and the subsequent evaluation of integrals, we follow Refs. Landshoff and Polkinghorne (1969); Parra-Martinez et al. (2020) and employ special variables where all mass and momentum-transfer dependence factorizes from the integrals, leaving nontrivial functions of the single dimensionless boost parameter $\sigma={p_{1}\cdot p_{2}}/{(m_{1}\,m_{2})}={1}/{\sqrt{1-v^{2}}}$, or equivalently the relative velocity $v$.
In this Letter, we focus on the potential region, i.e., the subset of the soft region describing instantaneous interactions between massive particles, where loop momenta scale as $\ell=(\ell^{0},\bm{\ell})\sim|q|(v,\mathbf{1})$ with $v\ll 1$. Classical observables are then extracted from the radial action, which is related to the massive-scalar $2\to 2$ scattering amplitude by Bern et al. (2021),
@@
Computational Setup— We extract classical conservative 5PM observables from the classical scattering amplitude of massive particles at $\mathcal{O}(G^{5})$ in general relativity, i.e. at four-loop order. We construct the integrand by first obtaining the requisite tree amplitudes via the double-copy, sewing them into generalized cuts, and then expanding in the soft region through ${\cal O}(|\bm{q}|^{2})$. The resulting expressions are then mapped to a cut-integrand basis Bern et al. (2025b) to remove redundant labeling and subsequently merged into a soft-expanded integrand by picking each term exactly once while verifying cut consistency.
@@
The IBP reduction is challenging due to the proliferation of irreducible numerators, doubled propagators, intricate topologies, and the large set of master integrals. We carry out the reductions with FIRE7 Smirnov and Zeng (2025), leveraging improved finite-field performance von Manteuffel and Schabinger (2015); Peraro (2016). In addition, we refine the choice of “seed integrals,” extending our previous strategy Bern et al. (2024b) and related developments Driesse et al. (2024); Guan et al. (2025). (see Kira 3 Lange et al. (2025) for a public implementation, Ref. Brunello:2025gpf for a recent application in the post-Newtonian expansion, and Refs. Larsen and Zhang (2016); Wu et al. (2023); Abreu et al. (2017, 2018) for analogous seeding schemes in earlier non-Laporta approaches.)
@@
As in Ref. Bern et al. (2025a), the differential equations determine the complete PM amplitude. Here, instead of closed-form expressions in terms of special functions, in the potential region, we exploit that the master integrals are meromorphic in $\bar{\sigma}\equiv\sigma-1$ near $\bar{\sigma}=0$ and thus admit a local Laurent expansion about the static limit. Solving the differential equations by the Frobenius method yields a series solution in principle to any desired precision,
@@
with coefficients $\vec{c}_{n}(\epsilon)$ that are rational functions of the dimensional regulator $\epsilon$. The lower bound $n_{0}$ is determined by the leading velocity scaling in the static limit, while $n_{\max}$ is chosen to reach the target accuracy, here $\mathcal{O}(\bar{\sigma}^{32})$. Expanding $\hat{M}(\sigma,\epsilon)$ around $\bar{\sigma}=0$, we solve the system recursively in $\bar{\sigma}$: at each order, the unknowns follow from linear systems whose $\epsilon$ dependence is reconstructed Wang (1981); Wang et al. (1982) from numerical evaluations over prime fields. This approach builds on earlier applications in collider physics Mistlberger (2018); Moriello (2020); Pozzorini and Remiddi (2006); Hidding (2021). Finally, after planarizing the most challenging 2SF classical contributions Bern et al. (2024b, 2025a) using the support of cut matter lines Saotome and Akhoury (2013); Akhoury et al. (2021), we obtain 371 boundary conditions, fixed from the static limit of scalar integrals that contain both classically singular and regular terms. Of these, only 95 enter the genuinely classical observables.
@@
5PM Results— Once expressed in terms of boundary integrals, the results organize themselves into well-defined eikonal sums Saotome and Akhoury (2013); Akhoury et al. (2021). The diagrams within a given eikonal sum are related by permuting vertices on matter lines, resulting in a sum over eikonal propagators that gives an on-shell $\delta$-function, forcing them to be cut. Fig. 2 gives the planar representatives defining the eikonal sums relevant for the classical amplitude, with the remaining ones given by permuting the attachment point of the potential-region graviton lines, represented by dashed lines, on the matter lines. See Refs. Bern et al. (2024b, 2025a) for further details on eikonal sums in the 5PM amplitudes.
@@
We solved the system of differential equations through $\mathcal{O}(\bar{\sigma}^{32})$; however, spurious singularities limit the result to the first 22 orders in the velocity expansion. Including the overall factor of $G^{5}$, this corresponds to terms through 26PN order.
@@
where for the sake of brevity we wrote explicitly only terms through 7PN order; further terms through $\mathcal{O}(\bar{\sigma}^{22})$ (26PN order) are included in the ancillary file GR_potential_radial_action.m. Such divergent terms in the potential-graviton contribution to the radial action are expected to be related to the odd-in-velocity part of the energy loss at lower (in this case 4PM) order Bini and Damour (2017); Bini et al. (2020); Blanchet et al. (2020),
@@
Terms through $\mathcal{O}(\bar{\sigma}^{22})$ are included in the ancillary file GR_potential_radial_action.m. The ancillary file also contains the exact analytic expressions for the 0SF and 1SF contributions. In these expressions, elliptic contributions cancel, as observed earlier in Ref. Driesse et al. (2024). The remaining functions are expressible in terms of manifestly real cyclotomic harmonic polylogarithms (CPL) Ablinger:2011te; Bern et al. (2024b) (c.f. also Ref. Ablinger:2021fnc for real representations of functions beyond CPLs). We note here the presence of terms proportional to $\pi^{2}$, which are absent from the analogous results in ${\cal N}=8$ supergravity Bern et al. (2025a). In our exact 1SF potential radial action, as well as in the 1SF conservative results of Ref. Driesse et al. (2024), such terms originate in (multiple) polylogarithmic functions. This is unlike the 4PM case, where $\pi^{2}$ terms tag only elliptic integrals. It will be interesting to determine the special functions that come with $\pi^{2}$ factors at 2SF order.
@@
Checks— Our result passes several nontrivial checks. First, as pointed out above, certain spurious higher-order poles in the dimensional regularization parameter $\epsilon$ cancel in the radial action and the remaining divergence matches the energy loss at 4PM order. At 0SF, we reproduce the exact probe limit result presented in e.g. Ref. Kol et al. (2022). Following the discussion in Ref. Bern et al. (2019a), the leading in velocity terms of the scattering angle are predicted by lower-order iterations. These velocity iterations arise from the integral representation of the scattering angle in terms of the radial momentum, see Eq. (2), and they determine the angle at orders $\bar{\sigma}^{-j}$, $j=2,3,4,5$,
@@
where $A=\frac{55}{6}{-}16\ln 2$. We checked using the 4PM angle $\chi_{4}$ Bern et al. (2021) that our results in Eq. (17) agree with this prediction. Note that the ${1}/{\epsilon}$ divergence of the 2SF angle is partially checked by the velocity iteration consistency at order $\bar{\sigma}^{-2}$ and the presence of a similar divergence of the potential-region 4PM angle $\chi_{4}$. Notably, a combination of all these checks targets all classical eikonal sums, reinforcing the correctness of our result. We have compared with the scattering angle that follows from the PN Hamiltonian of Refs. Blümlein et al. (2020, 2021, 2022) and found agreement.
Discussion— We include in Fig. 3 plots of the finite part of the radial action that illustrate some of their features. The left panel illustrates the relative importance of the different self-force components in the finite part of the radial action as a function of velocity. The right panel indicates that for $v\gtrsim.8$, higher-order terms and resummations will improve convergence. Once the radiation-mode contributions are incorporated, it will be essential to obtain accurate representations of the series solutions across the entire physical region. Several complementary strategies are available. As emphasized in Ref. Bern et al. (2025a), the choice of expansion variables plays a critical role in convergence. The conventional PN expansion in the velocity $v$ is not optimal, as it can introduce limitations from singularities outside the physical domain. A more suitable choice of expansion variable is $x$, related to the boost variable by $\sigma=(1+x^{2})/(2x)$, which leads to better convergence properties in the physical region $x\in[0,1]$. This strategy is incorporated in Fig. 3. See Ref. Ruf:Amplitudes23 for a discussion of spurious singularities of the solution to the differential equations.
@@
Conclusions— In this Letter, we computed the complete potential-mode contributions to the amplitude and conservative scattering angle at 5PM order. Our calculations show that the computationally demanding 2SF component can be brought under systematic control. We employed state-of-the-art Feynman integration tools, as published by two of the authors Smirnov and Zeng (2025), along with various private enhancements. Our framework closely parallels that employed in the corresponding analysis of maximal supergravity Bern et al. (2025a), augmented by substantial improvements in IBP seeding strategies, which should likewise transfer to multi-loop collider-physics applications.
After obtaining the four-scalar integrand and performing the IBP reduction, the differential equations and boundary integrals completely fix the PM amplitude and scattering angle. We explicitly solved the differential equations for all master integrals as a series expansion about the static limit $\bar{\sigma}\to 0$. The resulting solution determines the classical amplitude (radial action) at $\mathcal{O}(G^{5})$ through 26PN order. Extending the series to substantially higher orders is, of course, doable using the differential equation Eq. (7) and might be useful once radiation contributions are included. It would be interesting to promote the series solution to an exact analytic result, with full control over generalized polylogarithms, elliptic integrals, and Calabi–Yau integrals.
To complete the 5PM dynamics, radiation modes must also be included. While observables are independent of the precise split between conservative and dissipative dynamics, a definite scheme is required for comparison with other approaches Wheeler and Feynman (1949); Damour and Esposito-Farese (1996); Damour (2016); Barack and Pound (2019); Driesse et al. (2024). For waveform modeling, the PM results can be embedded in the EOB framework Khalil et al. (2022); Buonanno et al. (2024), potentially supplemented by resummations exploiting the singularity structure of observables Damour and Rettegno (2023); Long et al. (2024) or information from the high energy limit Alessio et al. (2025). Finally, extending the results to bound orbits requires an analytic continuation using appropriate prescriptions, as discussed in Refs. Khalil et al. (2022); Bini and Damour (2024); Dlapa et al. (2024, 2025). Such continuations are nontrivial at 4PM and beyond due to the nonlocal-in-time contributions from the tail effect. Once the complete 5PM results are available, it would be very interesting to compare them with high-precision numerical relativity Long et al. (2025) or self-force studies Barack et al. (2023) to assess the magnitude of the effects of time nonlocality (which may be small cf. e.g. Ref. Buonanno et al. (2024)) and the size of the still-to-be-calculated 6PM contributions as well as to identify improved resummations of the PM series Long et al. (2024).